QUESTION4 Which of the following t-scores is correct given the following information? A researcher is comparing the effectiveness of two sets of instructions for building a windmill. Two groups of individuals are randomly selected, and times to completion for Group 1 (n-4) were: 2.5, 3.0, 4.5., and 2.0 hours. Times to completion for Group 2 (n-4) were: 1.0, 2.0, 3.0, and 2.0 hours. O t-0.65 O t- 1.48 O t. 1.65 Ot 2.48
Answer #1

The answer for the above-mentioned question is

t=\pm2. 48

Hear no of individuals in Group1=4 and Group 2=4

Therefore d.f=(n1-1)+(n2-1)=(4-1)+(4-1)=6

d.f=6

Now using statistics t table with d.f =6 and alpha=0.05 ,  

t=\pm 2.48